2. Find b 5.0 35° 3. Find a & θ 9.5 84° 1.0 4. Find φ, θ, & c 4.0 3.0
Answer #1

2)
  For the triangle, we have
     5 cos 35
  b= 5 × cos 35° 4.0958

3)
From the definition of the sine, we have
     Cl sin 84° 9.5
  a 9.5 × sin 84° 9.448
And for the right angle triangle, we have
   θ 84°-90°
θ 900-84
Rightarrow heta=6^circ

4)
From the definition of the tangent, we have
     anphi=rac{3}{4}
  tan -1
() = 36.87

And for the right angle triangle, we have
   heta+phi=90^circ
And so,
   Rightarrow heta=90^circ-phi
  Rightarrow heta=90^circ-36.87^circ
  θ= 53.13

And from Pythagorus theorem, we have
   c^2=3^2+4^2
  Rightarrow c^2=9+16
C2-25
  Rightarrow c=sqrt{25}
  Rightarrow c=5